# Automating Governing Knowledge Commons and Contextual Integrity (GKC-CI) Privacy Policy Annotations with Large Language Models

### v2

@@ 4.3.1. Baseline Models
We consider two baselines for comparison against our fine-tuned LLMs: 1) a recurrent neural network (RNN) representing classical NLP approaches and 2) prompted non-fine-tuned LLMs.
For the RNN, we define BOS and EOS tokens, which is standard practice for these models (Robertson, 2024). We train the RNN in the exact same way as our fine-tuned LLMs, described in Sections 4.4.
For the prompted non-fine-tuned LLMs, we use GPT-4, GPT-4 Turbo, and GPT 3.5 Turbo. We use a fixed prompt format where we clearly delineate instructions from other text by placing instructions within brackets, thereby providing a structured framework that helps the model distinguish between the task directives and supplementary information. We also employ n-shot learning, a common prompting technique which has been shown to improve performance as n, the number of examples, increases  (Zamfirescu-Pereira et al., 2023; Brown et al., 2020; Gao et al., 2021). We use n={1, 3, 5} where the examples are chosen randomly from the training set.
@@ 4.3.2. Fine-Tuned Models
We consider five model families of diverse size and architecture: Flan-T5, GPT-2, Llama 2, GPT-3, GPT-4, and GPT-3.5 Turbo (Chung et al., 2022; Radford et al., 2019; Touvron et al., 2023b; Brown et al., 2020). Their properties are summarized in Table  2. In selecting models from these families, we wanted to choose from a wide range of high-performing or particularly usable LLMs. Approximately half of the models are open-source, while the GPT-3 models are proprietary. While the sizes of the GPT-3 models are not publicly released, GPT-3 and GPT-3.5 are likely the largest models we consider (their exact sizes are not public). We omit GPT-4 from our analysis because at the time of writing, the fine-tuning API is experimental. We also omit Llama2’s chat version from our analysis because, at the time of writing, it does not appear that Meta intended it to be fine-tuned further based on its GitHub repository (Facebook Research, 2023).
@@
We now give a quick summary of how the various architectural features of the models we trained may potentially impact performance on the annotation task. First, we vary model size. Specifically, we consider models both in their “base” or default size [^5] When loading from the HuggingFace model hub as well as a larger size, if permissible by our hardware [^6] We employ an NVIDIA A100 for training, which we believe to be a reasonable baseline for the amount of compute other researchers will have access to.. We consider model size because the number of parameters in a model plays a large role in its performance (Hoffmann et al., 2022; Kaplan et al., 2020; Touvron et al., 2023a; Chung et al., 2022). However, smaller models may approach larger models’ performance if they are given more (pre)training data. We consequently include the Llama2 model to serve as an open-source proxy for larger models’ (e.g. GPT-3) performance  (Touvron et al., 2023a, b).
@@
Most newer LLMs employ a “decoder-only” architecture, a deviance of the original design of the Transformer (Vaswani et al., 2017). Such models are referred to as encoder-decoder models, while models lacking an encoder block are decoder-only. Decoder only models can only ”look” at the preceding left tokens to determine what should be generated next, and tend to excel at creative or free-form generation  (Holtzman et al., 2020). In contrast, Encoder-Decoder models look at the entire sequence to determine what should be generated. Encoder-decoder models, like Flan-T5 or BERT, tend to perform well on tasks where the output is highly scoped by the input (Holtzman et al., 2020; Chung et al., 2022; Devlin et al., 2018). We include both Decoder-only and Encoder-Decoder models in our analysis because it is unclear what advantage, if any, an encoder-decoder model may have when annotating policies.
@@
There exist a number of training paradigms which result in a model becoming more aligned to human intent. By alignment, we specifically say that a model is aligned if it produces outputs which are consistent with its human operator’s desires (assuming the model is capable of producing those outputs.)  (Chen et al., 2021; Ouyang et al., 2022). Because alignment broadly reflects a models’ ability to output text consistent with input tasks, we hypothesize that models which have undergone specific alignment training may perform better. We specifically look at two alignment mechanisms: Instruction-Finetuning and Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022; Wei et al., 2022). We include models in our analysis which are either confirmed or likely to have been trained according to these techniques.
@@
Finally, we note that some models have been released as chat-models. This is relevant because 1) such models can be prompted in a different format from other non-chat models, and 2) these models are generally newer. While we are loath to conflate model age with performance, it is at least in the case of GPT and Llama that newer releases tend to eclipse older releases (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a, b). Thus, along with reasons explained in section 4.4, we include the fine-tunable chat model GPT-3.5 Turbo in our analysis.
@@ 4.4. Model Training
All 16 manually annotated privacy policies used for model training and testing underwent processing as outlined in section 4.2. We randomly reserved 70% of the manual annotations to constitute our training data (21,588 examples), while the other 30% (9252 examples) are reserved as testing examples. We specifically use low-rank adaption (LoRA), a type of parameter efficient fine-tuning (PEFT), as our training method instead of traditional fine-tuning to ensure our open-source models are being trained in a way which is consistent with proprietary models (Hu et al., 2021). PEFT methods in general encompass methods of training some of a model’s parameters, but not all (i.e., being ”parameter efficient” when training). LoRA describes a method of being parameter efficient when training by freezing a model’s weights and inserting trainable rank-decomposition matrices into the model’s architecture at each layer (typically, the attention layer). Then, the low-rank matrices are ”trained” during fine-tuning, while the base model remains frozen. We use LoRA because OpenAI’s business model suggests that LoRA is being employed in the place of traditional fine-tuning, as low rank matrices are very cheap to store and entire models are typically large and expensive. [^7] We make this claim because OpenAI allows for “fine-tuning” through their API. Traditional fine-tuning would require making full copies of the model (e.g., GPT-3.5) for each user. Doing so would result in terabytes of space being allocated per user due to the size of OpenAI’s models. Additionally, it has been observed that previously fine-tuned models under OpenAI may change in performance without warning, as OpenAI routinely updates their models. This behavior would not be observed if traditional fine-tuning were occurring because each user would have their own discrete copy of the model.
Thus, to ensure that all our model comparisons are apples-to-apples comparisons, we ensure that all models are trained using LoRA. We keep the training parameters constant between all open-source models. We make an exception for the learning rate of Flan-T5, as the model’s documentation recommends a slightly higher learning rate than the other models. This higher learning rate could cause the model to converge faster than others.
In training, we perform a number of ablation experiments, changing the following in a full-factorial manner for the open-source models:
@@
    Training Epochs, where we vary the number of epochs the model trains for. Training a model for a single epoch means that the LLM sees every example in the training set once. Training for n epochs means that the LLM sees every example in the training set n times. Varying the number of training epochs thus influences the quantity of training tokens the model is exposed to, as well as the number of repeated examples the model is exposed to.
@@
    Example formatting, where some examples are prepended and appended using the model’s BOS, EOS tokens. BOS (Beginning of Sentence) and EOS (End of Sentence) are typically defined and used during LLM pre-training to semantically capture the start and end of a training document. We hypothesize that the inclusion or exclusion of these tokens during fine-tuning may have a subtle effect on the model’s performance.
We note that Open-AI’s proprietary models do not offer the same number of training options as the open-source models, and are ultimately unable to perform the example formatting experiment with Open-AI’s models. We ultimately benchmark 50 models distributed between Open-AI’s and our open-source models.
@@ 5.1. Metrics
We begin a discussion of performance by clearly defining our performance metrics. Specifically, our main metric is accuracy. That is, we consider the model’s output to be correct or accurate if it is an exact string match of the ground-truth. We do not directly consider precision and recall because they are less intuitive for this task. We envision that downstream users are likely to be particularly interested in the performance per GKC-CI parameter. As such, we focus on per-parameter accuracy.
As we want to capture nuanced model errors, we categorize each model output into exactly one of four possible results: perfect match, superset match, match error, or identification error. These categories are defined as follows:
@@
    Perfect Match indicates the model’s annotation is an exact string match with that of the human annotator. Only the Perfect Match represents a correct annotation.
@@
    Superset Match indicates the model’s annotation contains all the words of the human annotator. However, the model may have highlighted additional information or may have included information which does not appear as a contiguous lump of text in the policy.
@@
    Match Error captures instances in which the model agreed that a certain parameter was present, but did not identify the “correct” annotation. This can include completions that are flat-out incorrect, completions that don’t identify the correct number of instances of a parameter in the input, and completions that have identified a proper subset ($\subset$) of the correct words.
@@ 5.2. Baseline Results
Our baseline RNN model had a test set accuracy of only 6% (Figure 1). While the RNN appropriately handled negative examples, it struggled with positive examples. Specifically, the RNN failed to accurately annotate any positive examples, while all LLMs other than GPT-2 were able to successfully annotate at least one positive example.
The prompted non-fine-tuned LLMs performed even worse, with the best such model having a test set accuracy of only 2% (Appendix D). While more effective prompting strategies are likely to exist, we believe that this finding provides evidence that LLMs should not be applied off-the-shelf for GKC-CI annotation as of now.
@@ 5.3. Fine-Tuned Results
We benchmark each of our 50 models on each of the 9252 sentences in our test set. The results of our benchmark are shown in Figure 1. We find that OpenAI’s proprietary models perform significantly better than any of the other open-source models we test. Additionally, no open-source model that we considered had performance high enough to be considered even a poor substitute. Examination of the open-source model’s outputs indicates a lack of substantial training convergence within 10 epochs for all models. Graphs of the open source models’ performance at 1, 5, and 10 epochs are included in Appendix E. We comment that each of the open-source models took ¡12 hours to train, but that training time is extremely sensitive to specific hardware configurations. We hypothesize that given enough compute, the open-source models may eventually reach the performance of Open AI’s proprietary models. We briefly summarize some notable lessons from working with the open-source models, then turn our discussion to the OpenAI models.
Figure 1. Test set performance of the top-performing models variants, including the RNN, with $\leq$ 10 epoch of training. GPT3,5_TPE refers to the prompt-engineered version of GPT-3.5 Turbo, GPT3,5_TG refers the generic GPT-3.5 Turbo model, and GPT3,5_t2s refers to the joint performance of the GPT-3.5 Turbo, 2-Step models.
@@ 5.3.1. Open-Source Takeaways
Firstly, we note that model size does play a significant role in model ability in relation to our annotation task. GPT-2 and Flan-T5 both failed utterly at their smaller model sizes, while their XL and Large variants performed $\sim$ 3% to 15% better. We additionally note that absolute model size in terms of parameters does not appear to be a consistent indicator of performance across model families, at least at the small sizes we consider. Llama2’s 7B variety appears to perform similarly to GPT2’s XL variety, despite being over four times the size. We mention this to caution resource-constrained researchers: larger may not always mean better when performing cross-model family comparisons. Additionally, those models which were aligned (Flan-T5 and all GPT3.5s) appeared to perform the best. We recommend researchers prioritize smaller, aligned models.
Secondly, we observe that models appear to be vulnerable to the inclusion or exclusion of BOS, EOS tokens. Namely, we observe small performance differences between all open-source models, but most strikingly in the Llama2 model family. We believe this is significant because it suggests that opaque defaults have observable effects on performance. Further, those effects do not appear to be consistent across model families. Specifically, as of the time of writing, HuggingFace’s tokenizers all have BOS, EOS tokens internally defined, but each model has a different default behavior when it comes to including or excluding BOS, EOS tokens. We urge other researchers to be particularly careful of library and model defaults as they could be a potential confound in model performance.
Figure 2. Per GKC-CI parameter performance for our best performing model, GPT 3.5TPE_25ep.
@@ 5.3.2. OpenAI Models
Appendix F summarizes model ability on positive examples, which are those examples that contain a GKC-CI parameter. We observe that models vary substantially in their per-parameter performance.
@@
All three models are variants of GPT-3.5 Turbo. GPT3.5-Turbo, two step 10 epochs refers to a two model system where the first model determines if the text contains a GKC-CI parameter and the second model identifies the parameter. For this model, an example is a perfect match if and only if the first model and second model classifies the prompt correctly. This model performed the worst out of the three at 30% correct positive examples and 97% correct negative examples. While the model demonstrates commendable success in identifying negative examples, its limited capacity to accurately identify positive examples significantly undermines the overall efficacy of its performance.
@@
GPT3.5-Turbo, Generic 10 epochs and GPT3.5-Turbo, Prompt Engineered 10 epochs performed similarly well with GPT3.5-Turbo, Generic 10 epochs at 73% correct positive examples and 88% correct negative examples and GPT3.5-Turbo, Prompt Engineered 10 epochs (hereafter, GPT 3.5TPE_10ep) at a slightly better 75% correct positive examples and 89% correct negative examples.
@@
Due to this strong performance, we further examine GPT 3.5TPE. We speculate that the reason why GPT 3.5TPE_10ep performs better than GPT 3.5TP_10ep is due to the added context from prompt engineering. Specifically, prompt engineering may provide enough context for GPT 3.5TPE_10ep such that the model is able to relate the annotation task to information it has been trained on. This phenomenon has been more broadly observed in the NLP space (Scao and Rush, 2021; Webson and Pavlick, 2022).
Like the open-source models, we similarly investigate how the number of training epochs effects the performance of this model. The results of this are shown in Figure  4. We observe that GPT 3.5TPE starts to have diminishing returns after 25 epochs worth of training. We hypothesize that at 50 epochs, it is likely that the model has overfit. We further investigate the effect that training data quantity has in Figure 4. We note that the model starts to have diminishing returns in terms of performance per data quantity at the 75% mark, and that 50% of our training data appears to be the necessary minimum for our model to begin to converge. Because we observe no detriment to using the full quantity of our training data, we do not restrict ourselves to a subset of our training data.
Finally, to ensure that our model’s performance isn’t critically hindered by being trained exclusively on sentences, we tested a fine-tuned version of GPT 3.5TPE, our most successful model, using paragraphs instead of sentences. The goal was to determine whether providing more context in the input would improve performance. Despite subjecting this model to the same epoch ablation experiments as our top-performing model, it only achieved a maximum accuracy of 67.75%. For a detailed comparison of the performance between both epoch ablation experiments, please refer to Appendix C.
To summarize, we observe that the model benefits from the full training data and that the model may be at risk of overfitting after 25 epochs. We consequently choose our top performing model to use the full training data, and train it for 25 epochs, which took 3 hours. We specifically call our top model GPT 3.5TPE_25ep, and its performance is shown in Figure  2. To better understand the strengths and weaknesses of this model, we perform qualitative error analysis on the model’s errors.
@@
Figure 3. GPT 3.5TPE’s performance on the test set at 1, 5, 10, 25, and 50 epochs. Perfect Matches and Superset Matches were considered to be ”correct”.
@@
Figure 4. GPT 3.5TPE_25ep’s accuracy on the test set as training data increases. Perfect Matches were ”correct.”

### v3

@@ 4.3.1. Baseline Models
We considered two baselines for comparison against our fine-tuned LLMs: 1) a recurrent neural network (RNN) representing classical NLP approaches and 2) prompted non-fine-tuned LLMs.
For the RNN, we defined BOS and EOS tokens, which is standard practice for these models (Robertson, 2024). We then perform standard training for the RNN but keep the rest of the training parameters consistent with the LLMs as described in Section 4.4.
For the prompted non-fine-tuned LLMs, we used GPT-4, GPT-4 Turbo, and GPT-3.5 Turbo. We used a fixed prompt format where we clearly delineated instructions from other text by placing instructions within brackets, thereby providing a structured framework that helped the model distinguish between the task directives and supplementary information. We also employed n-shot learning, a common prompting technique that has been shown to improve performance as n, the number of examples, increases  (Zamfirescu-Pereira et al., 2023; Brown et al., 2020; Gao et al., 2021). We used n={1, 3, 5} where the examples were chosen randomly from the training set.
@@ 4.3.2. Fine-Tuned Models
We considered five model families of diverse size and architecture: Flan-T5, GPT-2, Llama2, GPT-3, GPT-3.5 Turbo, and GPT-4 (Chung et al., 2022; Radford et al., 2019; Touvron et al., 2023b; Brown et al., 2020). Their properties are summarized in Table  2. In selecting models from these families, we wanted to choose from a wide range of high-performing or particularly usable LLMs. We ultimately omitted GPT-4 from our analysis because at the time of writing, the fine-tuning API was experimental. We also omitted Llama2’s chat version from our analysis because, at the time of writing, it did not appear that Meta intended it to be fine-tuned further based on its documentation (Facebook Research, 2023). Of the models we selected, approximately half are open-source, while the GPT-3 and GPT-3.5 models are proprietary. While the exact sizes of the GPT models are not publicly released, they are likely the among the largest of the models we tested.
@@
We now give a quick summary of how the various architectural features of the models we trained may impact performance on the annotation task. First, we varied model size. Specifically, we considered models both in their “base” or default size [^5] When loading from the HuggingFace model hub as well as a larger size, if permissible by our hardware. [^6] We employed an NVIDIA A100 for training, which we believe to be a reasonable baseline for the amount of compute other researchers will have available. We considered model size because the number of parameters in a model plays a large role in its performance (Hoffmann et al., 2022; Kaplan et al., 2020; Touvron et al., 2023a; Chung et al., 2022). However, smaller models may approach larger models’ performance if they are given more (pre)training data. We consequently included the Llama2 model to serve as an open-source proxy for larger models’ (e.g., GPT-3) performance  (Touvron et al., 2023a, b).
@@
Most newer LLMs employ a “decoder-only” architecture, a change from the original design of the Transformer (Vaswani et al., 2017). Such models are referred to as encoder-decoder models, while models lacking an encoder block are decoder-only. Decoder-only models can only “look” at the preceding tokens to determine what should be generated next and tend to excel at creative or free-form generation  (Holtzman et al., 2020). In contrast, encoder-decoder models look at the entire input sequence to determine what should be generated. Encoder-decoder models, like Flan-T5 or BERT, tend to perform well on tasks where the output is highly scoped by the input (Holtzman et al., 2020; Chung et al., 2022; Devlin et al., 2018). We included both decoder-only and encoder-decoder models in our analysis because it was unclear what advantage, if any, an encoder-decoder model might have when annotating privacy policies.
@@
There exist a number of training paradigms which result in a model becoming more aligned to human intent. We specifically mean that a model is aligned if it produces outputs which are consistent with its human operator’s desires (assuming the model is capable of producing those outputs) (Chen et al., 2021; Ouyang et al., 2022). Because alignment broadly reflects a model’s ability to output text consistent with input tasks, we hypothesized that models which have undergone specific alignment training might perform better. We considered two alignment mechanisms: Instruction-Finetuning and Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022; Wei et al., 2022). We included models in our analysis which are either confirmed or likely to have been trained according to these techniques.
@@
Finally, we noted that some models have been released as chat models. This is relevant because 1) such models can be prompted in a different format from other non-chat models (Section 4.2.1), and 2) these models are generally newer. While we are loath to conflate model age with performance, it is at least in the case of GPT and Llama that newer releases tend to eclipse older releases (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a, b). We therefore included the fine-tunable chat model GPT-3.5 Turbo in our analysis.
@@ 4.4. Model Training
We processed all 16 manually-annotated privacy policies into ground truth examples for model training and testing as described in Section 4.2. We randomly reserved 70% of the manual annotations to constitute our training data (21,588 examples), while the other 30% (9252 examples) were testing data.
We used low-rank adaptation (LoRA), a type of parameter efficient fine-tuning (PEFT), as our training method to ensure that the open-source models were trained in a way consistent with the proprietary models (Hu et al., 2021). PEFT describes a number of methods for training some, but not all, of a model’s parameters (i.e., being “parameter efficient” when training). LoRA is a PEFT method that involves freezing a model’s weights and inserting trainable rank decomposition matrices into the model’s architecture. These low-rank matrices are then “trained” during fine-tuning, while the base model remains frozen. OpenAI’s business model suggests that LoRA is being employed in the place of traditional fine-tuning, as low rank matrices are very cheap to store and entire models are typically large and expensive. [^7] We make this claim because OpenAI allows for “fine-tuning” through their API. Traditional fine-tuning would require making full copies of the model (e.g., GPT-3.5) for each user. Doing so would result in terabytes of space being allocated per user due to the size of OpenAI’s models. Additionally, it has been observed that previously fine-tuned OpenAI models may change in performance without warning, as OpenAI routinely updates their models. This behavior would not be observed if traditional fine-tuning were occurring because each user would have their own discrete copy of the model.
Thus, to ensure that all our model comparisons are fair, we trained all LLMs using LoRA. We kept the training parameters constant between all open-source models. We made an exception for the learning rate of Flan-T5, as the model’s documentation recommends a slightly higher learning rate than the other models. This higher learning rate could have caused the model to converge faster than others.
During training, we performed a number of ablation experiments, changing the following in a full-factorial manner for the open-source models:
@@
    Training Epochs. We varied the number of training epochs for the model. Training a model for a single epoch means that the model sees every example in the training set once. Training for n epochs means that the model sees every example in the training set n times. Varying the number of training epochs thus influences the quantity of training tokens to which the model is exposed, as well as the number of repeated examples to which the model is exposed.
@@
    Example formatting. We varied the examples’ formatting such that all training examples were prepended and appended with the model’s BOS, EOS tokens or not depending on the ablation condition. BOS (Beginning of Sentence) and EOS (End of Sentence) tokens are typically defined and used during LLM pre-training to semantically capture the start and end of a training document. We hypothesized that the inclusion or exclusion of these tokens during fine-tuning may have an effect on the model’s performance.
OpenAI’s proprietary models do not offer the same number of training options, so while were were able to experiment with different numbers of training epochs for OpenAI’s proprietary models, we were unable to perform the example formatting experiment described above with the proprietary models. We ultimately benchmarked 50 models distributed between open-source and OpenAI models.
@@ 5.1. Metrics
We begin our discussion of performance by clearly defining our metrics. Specifically, our main metric is accuracy. We consider a model’s output to be correct or accurate if it is an exact string match of the ground truth. We do not directly consider precision and recall because they are less intuitive for this task. We envision that downstream users are likely to be particularly interested in the performance per GKC-CI parameter. As such, we focus on per-parameter accuracy.
Since we want to capture nuanced model errors, we categorize each model output into exactly one of four possible results: perfect match, superset match, match error, or identification error. These categories are defined as follows:
@@
    Perfect Match indicates that the model’s annotation is an exact string match with that of the human annotator. Only a Perfect Match represents a correct annotation.
@@
    Superset Match indicates that the model’s annotation contains all the words of the human annotator. However, the model may have highlighted additional information or may have included information which does not appear as a contiguous sequence of text in the policy.
@@
    Match Error indicates that the model agreed that a certain parameter was present but did not identify the “correct” annotation. This can include completions that are flat-out incorrect, completions that don’t identify the correct number of instances of a parameter in the input, and completions that have identified a proper subset ($\subset$) of the correct words.
@@ 5.2. Baseline Results
Our baseline RNN model had a test set accuracy of only 6% (Figure 1). While the RNN handled negative examples well, it struggled with positive examples. Specifically, the RNN failed to accurately annotate any positive examples, while all LLMs other than GPT-2 were able to successfully annotate at least one positive example.
The prompted non-fine-tuned LLMs also performed poorly, with the best such model having a test set accuracy of $<$20% (Appendix D). While more effective prompting strategies are likely to exist, we believe that this finding provides evidence that LLMs should not be applied off-the-shelf for GKC-CI annotation as of now.
@@ 5.3. Fine-Tuned Results
We benchmarked each of our 50 models on each of the 9252 sentences in our test set. The results are shown in Figure 1. OpenAI’s proprietary models performed significantly better than any of the open-source models we tested. No open-source model that we tested had performance high enough to be considered even a poor substitute. Examination of the open-source models’ outputs indicates a lack of substantial training convergence within 10 epochs for all models. Graphs of the open source models’ performance at 1, 5, and 10 epochs are included in Appendix F. Each of the open-source models took ¡12 hours to train, but the training time is extremely sensitive to specific hardware configurations. We hypothesize that given enough compute, modern open-source models may eventually reach the performance of OpenAI’s proprietary models. We briefly summarize some notable lessons from working with the open-source models below. We then discuss the performance of the OpenAI models in greater detail.
Figure 1. Test set performance of the top-performing models variants, including the RNN, with $\leq$ 10 epoch of training. GPT3,5_TPE refers to the prompt-engineered version of GPT-3.5 Turbo, GPT3,5_TG refers the generic GPT-3.5 Turbo model, and GPT3,5_t2s refers to the joint performance of the GPT-3.5 Turbo, 2-Step models. Expanded model names in Appendix  E.
@@ 5.3.1. Open-Source Takeaways
First, we note that within a model family, size does play a significant role in a model’s ability to perform well on our annotation task. GPT-2 and Flan-T5 both failed utterly at their smaller model sizes, while their XL and Large variants performed $\approx$3to 15% better. We additionally note that absolute model size in terms of parameters across model families does not appear to be a consistent indicator of performance. Llama2’s 7B variety performed similarly to GPT-2’s XL variety, despite being over four times the size. We mention this to caution resource-constrained researchers: larger may not always mean better when comparing across model families. Additionally, those models which were aligned (Flan-T5 and all GPT-3.5 models) performed the best. We recommend compute-constrained researchers prioritize smaller, aligned models.
Second, we observe that models appear to be vulnerable to the inclusion or exclusion of BOS, EOS tokens. Namely, we observe small performance differences between all open-source models, but most strikingly in the Llama2 model family. We believe this is significant because it suggests that opaque defaults have observable effects on performance. Further, those effects do not appear to be consistent across model families. Specifically, as of the time of writing, HuggingFace’s tokenizers all have BOS, EOS tokens internally defined, but each model has a different default behavior when it comes to including or excluding BOS, EOS tokens. We urge other researchers to be particularly careful of library and model defaults as they could be a potential confound in model performance.
Figure 2. Performance per GKC-CI parameter for our best performing model, GPT 3.5TPE_25ep.
@@ 5.3.2. OpenAI Models
Appendix G summarizes model performance for three variants of GPT-3.5 Turbo on positive examples, which are those examples that contain a GKC-CI parameter. We observe that these models vary substantially in their per-parameter performance.
@@
GPT3.5-Turbo, two step 10 epochs refers to a two model system where the first model determines if the text contains a GKC-CI parameter and the second model identifies the parameter. For this model, an example is a perfect match if and only if the first model and the second model classify the prompt correctly. This model performed the worst of the three GPT-3.5 Turbo variants with 30% correct positive examples and 97% correct negative examples. While this model demonstrated success identifying negative examples, its limited capacity to accurately identify positive examples significantly undermined its overall performance.
@@
GPT3.5-Turbo, Generic 10 epochs (GPT 3.5TG_10ep) and GPT3.5-Turbo, Prompt Engineered 10 epochs (GPT 3.5TPE_10ep) performed similarly well to each other. GPT 3.5TG_10ep had 73% correct positive examples and 88% correct negative examples. GPT 3.5TPE_10ep had a slightly better 75% correct positive examples and 89% correct negative examples.
@@
We speculate that the reason why GPT 3.5TPE_10ep performed better than GPT 3.5TG_10ep is due to the added context from prompt engineering. Specifically, prompt engineering may provide enough context for GPT 3.5TPE_10ep such that the model is able to relate the annotation task to information on which it has been trained. This phenomenon has been more broadly observed in the NLP space (Scao and Rush, 2021; Webson and Pavlick, 2022).
@@
Figure 3. GPT 3.5TPE’s performance on the test set at 1, 5, 10, 25, and 50 epochs. Only Perfect Matches were considered to be “correct.”
@@
Figure 4. GPT 3.5TPE_25ep’s accuracy on the test set as training data increases. Only Perfect Matches were considered to be “correct.”
We next investigated how the number of training epochs effected the performance of GPT 3.5TPE (Figure 4). We observed that GPT 3.5TPE started to have diminishing returns after 25 epochs worth of training. This behavior is typical for any model trained using gradient descent with an optimizer since both the steps grow smaller as the model approaches a minima and the learning rate decays over time. We hypothesize that the model had likely overfit after 50 epochs. We further investigated the effect of training data quantity on GPT 3.5TPE (Figure 4). The model started to have diminishing returns in terms of performance per data quantity when using 75% of the training data, and 50% of the training data appeared to be the necessary minimum for the model to begin to converge. However, we observed no detriment to using all available training data.
To summarize, we observed that GPT 3.5TPE benefited from the full training dataset and that the model was at risk of overfitting after 25 epochs. We consequently chose GPT 3.5TPE as our best performing model and trained it for 25 epochs using the full training data, which took 3 hours. We call this top model GPT 3.5TPE_25ep, and its performance across GKC-CI parameters in the test set is shown in Figure 2. To better understand the strengths and weaknesses of this model, we performed a qualitative analysis of the model’s errors (Section 5.4).
Finally, to ensure that this model’s performance isn’t critically hindered by being trained exclusively on sentences, we fine-tuned a version of GPT 3.5TPE using paragraphs instead of sentences (details in Appendix  M). The goal was to determine whether providing more context in the input would improve performance. Despite subjecting this model to the same training epoch ablation experiments as our top-performing model, it only achieved a maximum accuracy of 67.75%, substantially lower than the 90.65% maximum accuracy achieved with sentence-based training. For a detailed comparison of the performance between these sentence-based and paragraph-based models, please refer to Appendix C.